Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
lunghini - 23:25 Friday 20 February 2026 (68725) Print this report
Operator Report - Afternoon shift

ITF found in COMMISSIONING Mode and LOCKING_ARMS_IR_ALS State, relock in progress after unlock and SAT activity with ITF DOWN.
Squeezing activity will resume when ITF will be in LOW_NOISE_3_ALIGNED.
All times are UTC.
15:02:38 ITF unlock from LOCKING_CARM_NULL_1F.
15:14:17 ITF unlock from LOCKING_CARM_NULL_1F.
15:25:13 ITF unlock from LOCKING_CARM_NULL_1F.
16:00 ITF in LOW_NOISE_3_ALIGNED.
16:22:22 ITF unlock from LOW_NOISE_3_ALIGNED.
16:57 ITF in LOW_NOISE_3_ALIGNED.
17:07:20 ITF unlock from LOW_NOISE_3_ALIGNED.
17:43 ITF in LOW_NOISE_3_ALIGNED.
20:32 Squeezing injection into ITF activity concluded and SQB1 Shutter (VPM > SQB1 > SQB1_Pico >  SQB1_Pico User Interface > ToITF_RetroRef) closed (Zendri, Vardaro, and Zambon, #68726).
20:46 Start of CEB Magnetic lines injections through ENV_MAIN Metratron node (operator: Lunghini, from remote: Tringali, Fiori).
ITF left in COMMISSIONING Mode and LOW_NOISE_3_ALIGNED State with AUTORELOCK_FAILSAFE STOP as asked by the experts.
CEB Environamental Magnetic Lines Injections are still running, and when ended ENV_MAIN Metatron node will go in DOWN State automatically.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×